The area between z=0 and z=2.86


1
Expert's answer
2022-04-12T05:52:57-0400

"P(0<Z<2.86)=\\
=P(Z<2.86)-P(Z<0)=\\
=0.9979-0.5000=0.4979 \text{ (from z-table).}"